OpenStudy (anonymous):

Find the arc length of a sector that is 125° and has a radius of 20 in. A. 13.9 B. 43.6 C. 872.2 D. 87.2

Directrix (directrix):

125/360 = x / [2 π r]

Directrix (directrix):

125/360 = x / [2 π r] 125/360 = x/ [2 π 20] 25/72 = x / 40π x = 125π / 9 where x represents the arc length

OpenStudy (anonymous):

L = r * A where A is in radians = 20 * 125 * pi --- 180 = 43.63 cm

Directrix (directrix):

B. 43.6

OpenStudy (amistre64):

C = angle * r; as in 2pi r